if AB is 4 cm longer than BC and the perimeter of ABCD is 68 cm, find the measures of AB,BC,CD and AD
Answer by Cromlix(4381)

Hi there,
I am assuming this is a rectangle.
Make the width = 'x' (BC)
Length = 4 + x (AB)
Perimeter of a rectangle =
2 x Width + 2 x Length
68 = 2(x) + 2(4 + x)
68 = 2x + 8 + 2x
Collect like terms
2x + 2x = 68 - 8
4x = 60
x = 15
AB = CD = 19 cm
BC = AD = 15 cm
Hope this helps :-)